# Learning the Intrinsic Dimensionality of Fermi-Pasta-Ulam-Tsingou Trajectories: A Nonlinear Approach using a Deep Autoencoder Model

### v1

@@ IV Results and Discussion
In Fig. 4 (top panel), some selected representative reconstruction error curves as functions of $m$. These curves correspond to $\beta=0.1,0.4,0.7,0.8,0.9,1$, and $\beta=1.1$. Additionally, a curve from PCA reconstruction error for $\beta=1.1$ is displayed. It has been rescaled by a factor of $100$ to enable visualization alongside the DAE curves, confirming the superior reconstruction performance of the autoencoder. All DAE reconstruction error curves decrease monotonically, except for the one corresponding to $\beta=0.1$, whose actual trend is not visible due to its small magnitude. For this reason, this curve is plotted separately in the bottom panel of Fig. 4 to improve visualization. All curves for which $\beta\leq 1$ exhibit an evident elbow point at $m=2$. Remarkably, this finding agrees with the approach based on multi-chart flows, recently proposed by Yu et al. [75]. By contrast, when $\beta=1.1$, the elbow point is shifted by unity, as confirmed by the Kneedle algorithm. Therefore, we can conclude that $m^{\ast}=2$ for $\beta\leq 1$, while $m^{\ast}=3$ for $\beta=1.1$.
We note in passing that one would then expect that a similar approach could be applied straightforwardly to reconstruction error curves at larger values of $\beta$. Nevertheless, this is not the case, as for increasing $\beta$ no evident elbows emerge, similarly to what happens in the linear approach. This fact is addressed in Sect. C, where the DAE and PCA curves correspond to $\beta=1.8$ (see Fig. 9). Consequently, alternative methods are required to ensure a reliable estimation of ID of trajectory data generated in the strongly nonlinear regime [77, 75, 7].
@@
Figure 6: Time evolution of the energy $E_{k}$ of the Fourier modes with $k=1,2,3,4,5$ in the FPUT $\beta$ model with $N=32$ oscillators. The initial condition corresponds to assigning the total energy $\mathcal{E}_{1}\approx 0.45$ to the first mode $E_{1}$ only.
@@
We can now compare the intrinsic dimensionalities inferred from the DAE with those previously obtained using PCA [52]. Since PCA admits several alternative heuristics to the Kneedle algorithm, we additionally consider the Kaiser criterion and the Participation Ratio, both of which are computed directly from the PCA eigenvalues (see Sects. B and B.1 for details on PCA and for the definitions of the heuristics, respectively).
@@
Fig. 6 clearly shows that PCA combined with KA (squares) and KC (triangles) substantially overestimates the intrinsic dimensionality, especially as $\beta$ increases, when compared with the DAE results (circles), which instead are consistent with the multi-chart flows approach. By contrast, PCA using PR agrees with the DAE estimates for $\beta\leq 0.3$ and for $\beta=1.1$, yielding $m^{\ast}=2$ and $m^{\ast}=3$, respectively. We argue that this agreement is purely accidental, since it is well known that PCA tends to overestimate the intrinsic dimensionality when data lie on a nonlinear manifold, as in the present case. A classical example is the synthetic Swiss roll dataset, for which PCA predicts $m^{\ast}=3$, whereas the data actually lie on a two-dimensional manifold ($m^{\ast}=2$[67, 72].
Additionally, within the linear approach the estimated intrinsic dimensionality remains unchanged as $\beta\rightarrow 1.1$, irrespective of the heuristic used, whereas the DAE predicts an increase to $m^{\ast}=3$.
@@
To understand this dimensionality variation, one needs to consider an important feature characterizing the $\beta$ model: the symmetry about its center for initially symmetric excitations [65, 59]. This means that if we initially excite a Fourier mode with an odd wave number, then the energy cannot spread to even-numbered modes. In the present case, the initial condition corresponds to assigning the total energy $\mathcal{E}_{1}\approx 0.45$ to the first mode $E_{1}$ only. As a result, the energy is observed to be shared by odd-numbered modes $E_{1}$, $E_{3}$, and $E_{5}$ only, in the regime of recurrence, i.e., when $\beta\leq 1$ (see, for instance, Fig. 8).
@@
In contrast, when $\beta=1.1$, even-numbered modes $E_{2}$ and $E_{4}$ start to be excited. Note that this happens only toward the end of the simulation; see Fig. 6, where $E_{2}$ and $E_{4}$ can acquire approximately a maximum of $6\%$ of the total energy. Therefore, we can relate the change in ID from $m^{\ast}=2$ to $m^{\ast}=3$ to the occurrence of SB in the $\beta$ model. We suspect that this variation in ID may imply a substantial difference in the geometric or topological features of the $2$-dimensional and $3$-dimensional manifolds on which the trajectories lie. Accordingly, this interesting finding could be properly investigated using other ML tools, for instance, those from topological and geometric data analysis [8, 10, 58, 56, 60, 33, 32].
One advantage of using autoencoders is the possibility of visualizing the data and thereby gaining meaningful insights into the dynamics. Accordingly, we compute a two-dimensional latent representation (or embedding) of the entire trajectory, consisting of $n_{s}=3\,000\,000$ data points, generated at very small nonlinearity, $\beta=0.1$  [^4] Here we used a smaller number $n_{s}$ of data points to reduce the computational burden.. In particular, we compare the two embeddings obtained when ReLU and Linear activation functions are employed in the bottleneck layer. Panels $(a)$ and $(b)$ of Fig. 7 show the resulting two-dimensional embeddings for the ReLU and Linear cases, respectively. In both cases, the embeddings define compact, ring-like regions with sharp, clearly delineated boundaries.
In particular, the embedding in panel $(b)$ confirms that the trajectory data is centered during the preprocessing step. This fact is not evident in the ReLU case, since the activation function clips negative values to zero. Note that the two activation functions yield small, comparable reconstruction errors, i.e., $J_{2}\approx 10^{-4}$, confirming that these DAE hyperparameters are essentially equivalent.
Note that the embeddings suggests that the motion is confined in a relatively small region of phase space. As a result, the orbits should lie on a small nonlinear manifold. In contrast, for larger values of $\beta$, for which the system reaches thermal equilibrium, the corresponding two-dimensional embeddings display irregular patterns, suggesting that the trajectory is now free to roam over a much larger region of phase space(unbounded motion). This dynamical behavior is illustrated in Fig. 10, where the data correspond to $\beta=3$.
@@ V Conclusion
In this work, we addressed the intrinsic dimensionality of high-dimensional trajectories in the FPUT chain using a nonlinear manifold learning approach. To this end, we employed a deep autoencoder model to analyze the trajectory data. Accordingly, we find that in the weakly nonlinear regime ($\beta\lesssim 1$) the orbits lie on a small two-dimensional nonlinear Riemannian manifold. In contrast, a comparison with a linear approach based on principal component analysis shows that it cannot capture the nonlinear structure of the data, thereby yielding questionable results. Additionally, we showed that PCA is unable to detect crucial dynamical changes, such as the symmetry breaking occurring at $\beta=1.1$. In this case, the autoencoder predicts an increase in the intrinsic dimension by unity, which may be related to significantly different geometric and topological features of the resulting three-dimensional nonlinear manifold.
In future work, it would be highly desirable to validate our findings using alternative approaches that do not rely on heuristics such as the detection of elbow points [77, 75, 7].
Finally, our work shows that a data-driven approach can deliver new and interesting insights into FPUT dynamics, since trajectories exhibit rich geometric features. Furthermore, the use of manifold learning methods, which rests on the manifold hypothesis, namely, that high-dimensional data lie near or on low-dimensional manifolds [15, 54], is supported by the applicability of the KAM theorem in the weakly nonlinear regime to $\beta$ model [44].
@@ Acknowledgements.
The author thank Søren Hauberg for advice on correctly training the deep autoencoder and Alberto Maspero for explaining the Cauchy problem for the FPUT system as a first-order differential system. Most of the code execution for this project was performed using Google Colaboratory. The author is also grateful to Google for supporting this research through the Tensor Research Cloud (TRC) program and for providing access to Cloud Tensor Processing Units (TPUs), which accelerated the machine learning experiments.
@@ Appendix B Principal Component Analysis
Let $\tilde{X}$ denote the $n_{s}\times n$ data matrix, after the data preprocessing, which make the variables scale-free each with zero mean, then the sample covariance matrix (or equivalently correlation matrix) $S$ reads [37]
$S=\frac{1}{n_{s}-1}\tilde{X}^{T}\tilde{X}\,$. (6)
The eigenvalue $\lambda_{i}$ of $S$ correspond to the explained (or preserved) variance along the $i$-th principal component axis. In present work, the sorted eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\cdots\lambda_{i}\geq\cdots\lambda_{n}$ are efficiently computed using the singular value decomposition [71, 70]. Accordingly to SVD, $\tilde{X}=WLV^{T}$ where $W$ and $V$ are two orthogonal matrices, and $L$ is a diagonal matrix [37, 28]. As a result, it is found that [37]
@@ B.1 Heuristics for Determining ID in PCA
An alternative way to obtain the ID from PCA’s eigenvalues is computing the Participation Ratio $D_{\rm PR}$ which estimates the number of dimensions along which the data is distributed [43, 63]. Accordingly, $D_{\rm PR}$ reads [63]
@@
Eq. 10 can be conventiently expressed in terms of the traces of $S$ (Eq. 6) and $S^{2}$, respectively. In such a case, one finds $D_{\rm PR}=\left(\operatorname{Tr}(S)\right)^{2}/\operatorname{Tr}(S^{2})$.
Finally, note that $D_{\rm PR}$ is a real number, so we shall round it to its nearest integer.
@@
Figure 9: DAE reconstruction error curve (circle) and PCA reconstruction error curve (triangle), rescaled by a factor of $200$, as functions of latent dimension $m$ for trajectory data corresponding to $\beta=1.8$.
Next, in Fig. 10, the two-dimensional DAE embedding of the entire trajectory with $n_{s}=3000000$ data points and $\beta=3$ is shown. In contrast to the embeddings obtained for small nonlinearities, i.e., $\beta=0.1$, there is no longer a well-defined pattern as observed previously. Furthermore, this embedding suggests that the motion is now unbounded, allowing the system to explore a much larger region of the phase space.
For sake of completeness, we report the intrinsic dimensionality of this trajectory data. It is found that PCA estimates $m^{\ast}=10$ (PR), $m^{\ast}=37$, and $m^{\ast}=37$ [52].

### v2

@@ IV Results and Discussion
In Fig. 5 (top panel), some selected representative reconstruction error curves as functions of $m$ are shown. These curves correspond to $\beta=0.1,0.4,0.7,0.8,0.9,1$, and $\beta=1.1$. Additionally, the curve from PCA reconstruction error for $\beta=1.1$ is displayed. The latter has been rescaled by a factor of $100$ to enable visualization alongside the DAE curves, confirming the superior reconstruction performance of the autoencoder. All DAE reconstruction error curves decrease monotonically, except for the one corresponding to $\beta=0.1$, whose actual trend is not visible due to its small magnitude. For this reason, this curve is plotted separately in the bottom panel of Fig. 5 to improve visualization. All curves for which $\beta\leq 1$ exhibit an evident elbow point at $m=2$. Remarkably, this finding agrees with the approach based on multi-chart flows recently proposed by Yu et al. [81]. By contrast, when $\beta=1.1$, the elbow point is shifted by unity, as confirmed by the Kneedle algorithm. Therefore, we can conclude that $m^{\ast}=2$ for $\beta\leq 1$, while $m^{\ast}=3$ for $\beta=1.1$.
@@
Figure 7: Time evolution of the energy $E_{k}$ of the Fourier modes with $k=1,2,3,4,5$ for the $\beta$ model with $N=32$ oscillators. The initial condition corresponds to assigning the total energy $\mathcal{E}_{1}\approx 0.45$ to the first mode $E_{1}$ only.
We note in passing that one would expect a similar approach, based on elbow detection, to be applicable straightforwardly to reconstruction error curves at much larger values of $\beta$, not considered in the present work. Nevertheless, this is not the case, since as $\beta$ increases no evident elbows emerge in the reconstruction error curve. This behaviour is analogous to that observed for PCA reconstruction error curves [57]. To illustrate the above observations, the DAE and PCA curves corresponding to $\beta=1.8$ are shown in Fig. 11 of Sec. C.
That being said, alternative methods will be necessary to ensure a reliable estimation of the ID of trajectory data generated in the strongly nonlinear regime of the FPUT $\beta$ model [83, 81, 7].
@@
We can now compare the intrinsic dimensionalities inferred from the DAE with those previously obtained using PCA [57]. Since PCA admits several alternative heuristics to the Kneedle algorithm, we additionally consider the Kaiser criterion (KC) [41] and the Participation Ratio (PR) [48], both of which are computed directly from the PCA eigenvalues (see Secs. B and B.1 for details on PCA and for the definitions of the heuristics, respectively).
@@
In Fig. 7 we compare the estimated IDs using the DAE model and KA (circles) with those from PCA together with PR (triangles), KA (squares), and KC (inverted triangles), respectively. First, we note that PCA with the KA and KC substantially overestimates the intrinsic dimensionality for $\beta\geq 0.3$, since in such cases $4\leq m^{\ast}\leq 6$. By contrast, PCA together with PR gives $m^{\ast}=2$ and $m^{\ast}=3$ when $\beta\leq 0.3$ and $\beta>0.3$, respectively. These results are in excellent agreement with those obtained from the DAE reconstruction error curves. Nevertheless, we argue that this agreement is purely coincidental, since it is well known that PCA tends to overestimate the intrinsic dimensionality when data have a nonlinear structure, as in the present case. A classical example is the synthetic Swiss roll dataset, for which PCA predicts $m^{\ast}=3$, whereas the data actually lie on a two-dimensional nonlinear manifold ($m^{\ast}=2$[72, 77]. Based on general considerations, PCA should overestimate IDs by at least unity when the data are nonlinear [83]. Consequently, we may conclude that in the present case the linear approach works properly only when $0.3\leq\beta\leq 1$.
@@
However, we note that there is another important difference between the DAE approach and the linear one. Indeed, according to the latter, the estimated ID remains unchanged as $\beta\leq 1.1$, irrespective of the heuristic used, whereas the DAE predicts an increase by unity, i.e., $m^{\ast}=3$. To understand this dimensionality variation, one needs to consider an important feature characterizing the $\beta$ model: the symmetry about its center for initially symmetric excitations [70, 64]. This means that if we initially excite a Fourier mode with an odd wave number, then the energy cannot spread to even-numbered modes. In the present case, the initial condition corresponds to assigning the total energy $\mathcal{E}_{1}\approx 0.45$ to the first mode $E_{1}$ only. As a result, the energy is expected to be shared only by the odd-numbered modes only, in the regime of recurrence. In our case, this occurs when $\beta\leq 1$, as illustrated for instance by the simulation for $\beta=1$ in Fig. 9. In contrast, when $\beta=1.1$, the even-numbered modes $E_{2}$ and $E_{4}$ are also involved in the system’s dynamics, as illustrated in Fig. 7. In such a case, the mode energies $E_{2}$ and $E_{4}$, shown as dotted curves in purple and cyan, respectively, acquire a small amount of energy towards the end of the simulation. However, this energy gain is not attributable to the symplectic integrator numerical error. Indeed, in Sec. A we show that the energy drop of odd modes $E_{1},E_{3}$ and $E_{5}$, which is approximately $6.55\%$ of total energy $\mathcal{E}_{1}$, quantitatively accounts for the energy transferred to even modes $E_{2}$ and $E_{4}$, corresponding approximately $6.02\%$ of $\mathcal{E}_{1}$ (see, in particular Fig. 10). This energy transfer from odd to even modes occurs slowly in a regime where the system has not yet reached thermal equilibrium and cannot be attributed to the characteristic oscillatory time dependence of the symplectic integrator error [16].
@@
That being said, we can relate the change in ID from $m^{\ast}=2$ to $m^{\ast}=3$ to the occurrence of SB in the $\beta$ model, whereby energy begins to flow from odd to even modes. Notably, this important dynamical feature could not be detected using PCA due to the nonlinear structure of the trajectory data. Accordingly, we suspect that the change in ID due to symmetry breaking may imply a substantial difference in the geometric or topological features of the $2$- and $3$-dimensional Riemannian manifolds on which the trajectories lie. This could be properly addressed using suitable tools from topological and geometric data analysis [9, 11, 63, 61, 65, 37, 36].
@@
Finally, one advantage of using autoencoders is the possibility of visualizing the data and thereby gaining meaningful insights into the dynamics. Accordingly, we compute a two-dimensional latent representation (or embedding) of the entire trajectory, consisting of $n_{s}=3\,000\,000$ data points, generated at very small nonlinearity $\beta=0.1$  [^5] Here we used a smaller number $n_{s}$ of data points to reduce the computational burden.. In particular, we compare the two embeddings obtained when ReLU and Linear activation functions are employed in the bottleneck layer. Panels (a) and (b) of Fig. 8 show the resulting two-dimensional embeddings for the ReLU and Linear cases, respectively. In both cases, the embeddings define compact, ring-like regions with sharp, clearly delineated boundaries.
Note that the embedding in panel (b) confirms that the trajectory data are centered during the preprocessing step. This is not evident in the ReLU case, since the activation function clips negative values to zero. In both cases the two activation functions yield small, comparable reconstruction errors $J_{2}\approx 10^{-4}$.
These two embeddings suggest that the system’s motion is confined to a relatively small region of phase space. As a result, the orbits should lie on a tiny two-dimensional nonlinear Riemannian manifold. In contrast, for larger values of $\beta$, for which the system reaches thermal equilibrium, the corresponding embeddings no longer show a compact ring-like structure, but irregular patterns, consistent with an ergodic exploration of a larger region of phase space. For instance, the characteristic irregular pattern emerging in the strongly nonlinear regime ($\beta=3$) is illustrated in Fig. 12, suggesting that the trajectory is free to explore a much larger region of phase space.
@@ V Conclusion
In this work, we addressed the intrinsic dimensionality of high-dimensional trajectories in the FPUT chain using a deep learning approach. To this end, we employed a deep autoencoder model to analyze the trajectory data. Accordingly, we find that in the weakly nonlinear regime ($\beta\lesssim 1$) the orbits lie on a small two-dimensional nonlinear Riemannian manifold. In contrast, a comparison with a linear approach based on principal component analysis shows that it cannot capture the nonlinear structure of the data, thereby yielding questionable results. Additionally, we showed that PCA is unable to detect crucial dynamical changes, such as the symmetry breaking occurring at $\beta=1.1$. In this case, the autoencoder predicts an increase in the intrinsic dimension by unity, which may be related to significantly different geometric and topological features of the resulting three-dimensional nonlinear manifold.
In future work, it would be highly desirable to validate our findings using alternative approaches that do not rely on heuristics, such as the detection of elbow points [83, 81, 7].
Finally, we conclude noting that the use of manifold learning methods, which rests on the manifold hypothesis, namely, that high-dimensional data lie near or on low-dimensional manifolds [17, 59], is strongly supported by the applicability of the KAM theorem in the weakly nonlinear regime to $\beta$ model [49].
@@ Acknowledgements.
The author thanks Søren Hauberg for advice on correctly training the deep autoencoder, Fabio Frohberg for helpful comments on the manuscript and for allowing the use of his JAX-based deep autoencoder script, and Alberto Maspero for explaining the Cauchy problem for the FPUT system as a first-order differential system. Most of the code execution for this project was performed using Google Colaboratory. The author is also grateful to Google for supporting this research through the Tensor Research Cloud (TRC) program and for providing access to Cloud Tensor Processing Units (TPUs), which accelerated the machine learning experiments.
@@ Appendix A FPUT Initial Condition
Next, we show that the energy transfer to modes $E_{2}$ and $E_{4}$ is a direct consequence of the symmetry-breaking feature of the $\beta$ model, rather than an artifact of symplectic integration errors. To this end, we define the following energy fractions: $\left(E_{1}+E_{3}+E_{5}\right)/\mathcal{E}_{1},\quad\left(E_{2}+E_{4}\right)/\mathcal{E}_{1}$, and $\left(\sum_{i=6}^{32}E_{i}\right)/\mathcal{E}_{1}$.
We then examine their time evolution during the system dynamics for $\beta=1.1$ (see Fig. 10). Toward the end of the simulation, the energy of the odd modes $E_{1}$, $E_{3}$, and $E_{5}$ (blue curve) decreases by approximately $6.55\%$, while the even modes $E_{2}$ and $E_{4}$ (brown curve) gain about $6.02\%$ of $\mathcal{E}_{1}$. The difference between these values is therefore small, about $0.53\%$, confirming that an energy redistribution between these modes is taking place.
We also note that the energy of the remaining modes with $k\geq 6$ (grey curve) oscillates within the range $0\%-20\%$, indicating that no systematic energy accumulation occurs in these modes. We therefore conclude that the observed symmetry breaking is not a fictitious effect arising from numerical integration errors.
Figure 10: Time-evolution of the energy fractions (in percentage) $\left(E_{1}+E_{3}+E_{5}\right)/\mathcal{E}_{1}$ (odd modes, blue curve), $\left(E_{2}+E_{4}\right)/\mathcal{E}_{1}$ (even modes, brown color) and $\left(\sum_{i=6}^{32}E_{i}\right)/\mathcal{E}_{1}$ (modes with $k\geq 6$, grey curve).
@@ Appendix B Principal Component Analysis
Let $X^{\ast}$ denote the $n_{s}\times n$ data matrix, after the data preprocessing, which make the variables scale-free each with zero mean, then the sample covariance matrix (or equivalently correlation matrix) $S$ reads [42]
$S=\frac{1}{n_{s}-1}X^{\ast^{T}}X^{\ast}\,.$ (6)
The eigenvalue $\lambda_{i}$ of $S$ correspond to the explained (or preserved) variance along the $i$-th principal component axis. In present work, the sorted eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\cdots\lambda_{i}\geq\cdots\lambda_{n}$ are efficiently computed using the singular value decomposition [76, 75]. Accordingly to SVD, $X^{\ast}=WLV^{T}$ where $W$ and $V$ are two orthogonal matrices, and $L$ is a diagonal matrix [42, 32]. As a result, it is found that [42]
@@ B.1 Heuristics for Determining ID in PCA
An alternative way to obtain the ID from PCA’s eigenvalues is computing the Participation Ratio $D_{\rm PR}$, which estimates the number of dimensions along which the data is distributed [48, 68]. Accordingly, this quantity reads [68]
@@
Eq. 10 can be conventiently expressed in terms of the traces of $S$ (Eq. 6) and $S^{2}$, respectively. In such a case, one finds $D_{\rm PR}=\left(\operatorname{Tr}(S)\right)^{2}/\operatorname{Tr}(S^{2})$. Since $D_{\rm PR}$ is as an estimate of the intrinsic dimensionality of the data, an inherently integer-valued quantity, we round it to the nearest integer.
@@
Figure 11: DAE reconstruction error curve (circle) and PCA reconstruction error curve (triangle), rescaled by a factor of $50$, as functions of latent dimension $m$ for trajectory data corresponding to $\beta=1.8$.
Next, in Fig. 12, the two-dimensional DAE embedding of the entire trajectory with $n_{s}=3\,000\,000$ data points and $\beta=3$ is shown. In contrast to the embeddings obtained for small nonlinearities, i.e., $\beta=0.1$, there is no longer a regular pattern. Additionally, the embedding suggests that the system is free to explore a much larger region of the phase space, consistently with the ergodic hypothesis.
We conclude reporting the intrinsic dimensionality of these trajectory data according to PCA [57]. Accordingly, it is found that $m^{\ast}$ ranges from $10$ (PR) to $37$ (KA, KC).